solve the polynomial equation by using the zero product principle.
x^3-27=0

Answer by stanbon(75887)

solve the polynomial equation by using the zero product principle.
x^3-27=0
----
Factor:
(x-3)(x^2+3x+27) = 0
---
Solve the quadratic factor:
x^2 + 3x + 27 = 0
x = [-3 +- sqrt(9-4*27)/2
---
x = [-3 +- sqrt(-99)]/2
--
x = [-3+-3sqrt(11)i]/2 (2 complex number solutions)
x = 3 (the Real Number solution)
====================================
Cheers,
Stan H.